OpenStudy (anonymous):

Write the following fraction as a decimal 2/3

OpenStudy (anonymous):

makes a lousy decimal; since \(\frac{1}{3}=0.3333...\) you know \(\frac{2}{3}=0.66666...\)

OpenStudy (anonymous):

sometimes written as \(\overline{.6}\)

OpenStudy (anonymous):

2/3 as a decimal is 0.6666666666666666 (it keeps going) but if you're doing a math problem that wants you to round it to the nearest hundredth, it would be 0.67 (because the 6 rounds to 7).

OpenStudy (anonymous):

2 apples divided among 3 people, if each person got 1/2 of an apple, that would be 0.5, and there's still another 1/2 left over. 0.5 / 3 is 0.1666. 0.5 + 0.1666 = 0.6666... no matter how many times you keep dividing it will keep coming out as 0.666, plus another 6, plus another six, etc.

OpenStudy (anonymous):

- yahoo answers;)

OpenStudy (anonymous):

This may not seem important, but since the denominator is 3, I thought I'd throw my chips in as well. Divide any number number by 9 and something cool happens:\[\frac{ 6 }{ 9} (i.e.2/3)=.6666666666....\]\[\frac{ 4 }{ 9}=.4444444....\]\[\frac{ 13 }{ 9 }=1\frac{ 4 }{ 9 }=1.444444....\]